# Explainable Concept Generation through Vision-Language Preference Learning

### v1

@@ 3.2 The Deep Reinforcement Learning Formulation
Explainable states: States that have a concept score $S_{c,m}\geq\eta$ for a user-defined threshold $\eta\in[0,1]$ for concept $c$ and class $m$ is defined as an explainable state.
@@
Proxy states: States that have a concept score $S_{c,m}<\eta$ for the threshold $\eta\in[0,1]$ for concept $c$ and class $m$ is defined as a proxy state.
@@
$Q^{*}(s,a)=\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[\xi_{t}r(s,a)+\gamma\max_{a^{\prime}\in A}Q_{\text{target}}(s^{\prime},a^{\prime})],with,$ (3)
  • Reward $r_{t}$: Reward $r_{t}$ is proportional to the TCAV score computed at state $s_{t}$ on action $a_{t}$, adjusted by a monotonically increasing scaling factor $\xi_{t}$. This factor is used to scale the reward over time $t$ Since the $g(.)$ is getting optimized at each timestep $t$. The scaling factor is updated as $\xi_{t+1}\leftarrow\min(1,\xi_{t}+k)$. Therefore, the expected cumulative adjusted reward is $R(\pi)=\mathbb{E}[\sum_{t=0}^{T}\xi_{t}\cdot r_{t}(s_{t},a_{t})]$.
3.3 Extracting the Action Space
Since a generative model can generate arbitrary images, if we provide good enough starting points, then the convergence to explainable states would be faster. Therefore, to automatically identify potential keywords, that becomes actions for the DQN, each test image used in TCAV, $X_{m}$, is cropped into nine patches. Note that, unlike segmentation based techniques that directly use image patches as concepts [5], because we use preference optimization in a latter stage, our algorithm merely considers information extracted from these patches as starting points. For instance, Fig. 2 shows the progression of images generated for the text prompt “zoo.” We then use a Visual Question Answering (VQA) model to ask about properties of each patch. The generated textual descriptions for all the patches undergo standard text preprocessing such as stop word removal, duplicate removal, etc. In its simplest form, we use the resulting keywords, $\mathcal{K}$ as the action space. In its simplest form, the DQN may pick only one keyword (e.g., “stripes”). Going beyond simple concepts, the algorithm can also select multiple keywords (e.g., “black and white” + “vertical” + “stripes”), in which case, the DQN can consider $N^{|\mathcal{K}|}$ permutations, though we limit this to $|\mathcal{K}|^{5}$ in our experiments.
Figure 2: Different levels of abstractions for the class “Tiger.” The generated image begins as a random zoo image and gradually transition to images with tiger features, ultimately transforming into a tiger itself.
@@ 3.4 Optimizing the States
At time $t$, the policy picks the keyword $k_{t}$, which is then used by the generative model, $g(k_{t};w_{t})$, with model weights $w$, to generate $2Z$ number of images. We randomly divide the generated images into two groups: $X_{c_{1},t}=\{x_{c_{1},t,i}\}_{i=1}^{Z}$ and $X_{c_{2},t}=\{x_{c_{2},t,i}\}_{i=1}^{Z}$. Let the TCAV scores of each group be $S_{c_{1},m,t}$ and $S_{c_{2},m,t}$. Since our objective is to find concepts that generate a higher TCAV score, concept images that have a higher score is preferred. Note that, unlike in the classical DPO setting with a human to rank, preference for our optimization comes from TCAV scores (e.g., $X_{c_{1},t}\succ X_{c_{2},t}$). If the generative model at time $t$ is not capable of generating concepts that are in an explainable state, $\max(S_{c_{1},m,t},S_{c_{2},m,t})\leq\eta$, we then perform preference update on SD’s weights. Following Low-Rank Adaptation (LoRA) [14], we only learn auxiliary weights $a$ and $b$ at each time step, and update the weights as $w_{t+1}\leftarrow w_{t}+\lambda ab$.
@@
As the DRL agent progresses over time, the states become more relevant as it approaches explainable states (Fig. 1), thus the same action yields increasing rewards over time. To accommodate this, with reference to the rewards defined in Section 3.2, we introduce a parameter $\xi$ which starts at 0.1 and incrementally rises up to 1 as the preference threshold, $\eta$, is approached. Different actions may result in different explainable states, reflecting various high-level concepts inherent to the NUT. Some actions might take longer to reach an explainable state. Also, it is possible for different actions to lead to the same explainable state. As the goal is to optimize all states to achieve a common target, DQN progressively improves action selection to expedite reaching these states. Thus, DRL becomes relevant as it optimizes over time to choose the actions that are most likely to reach an explainable state more efficiently.
3.5 The Effects of RLPO-based Concept Space Traversal
As depicted in Fig. 3, we can formalize the notion that a generative model can encompass concepts pre-defined by a human or retrieved through segmentation as follows.
@@ 3.5 The Effects of RLPO-based Concept Space Traversal
We now formalize what concepts the DQN has learned, with reference to Fig. 1.
@@
Property 1.
As $\xi$ escalates, the reward function proportionally amplifies, particularly enhancing the significance of outcomes near $t_{\eta}$.

### v2

@@ 3.3 Deep Reinforcement Learning Formulation
Explainable states: States that have a concept score $TS_{c,m}\geq\eta$ for a user-defined threshold $\eta\in[0,1]$ for concept $c$ and class $m$ is defined as an explainable state.
@@
Proxy states: States that have a concept score $TS_{c,m}<\eta$ for the threshold $\eta\in[0,1]$ for concept $c$ and class $m$ is defined as a proxy state.
@@
$Q^{*}(s,a)=\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[\xi_{t}r(s,a)+\gamma\max_{a^{\prime}\in A}Q_{\text{target}}(s^{\prime},a^{\prime})].$ (3)
@@ 3.4 Optimizing the States
At time $t$, the policy picks the seed prompt $k_{t}$, which is then used by the generative model, $g(k_{t};w_{t})$, with model weights $w$, to generate $2Z$ number of images. We randomly divide the generated images into two groups: $X_{c_{1},t}=\{x_{c_{1},t,i}\}_{i=1}^{Z}$ and $X_{c_{2},t}=\{x_{c_{2},t,i}\}_{i=1}^{Z}$. Let the TCAV scores of each group be $TS_{c_{1},m,t}$ and $TS_{c_{2},m,t}$. Since our objective is to find concepts that generate a higher TCAV score, concept images that have a higher score is preferred. Note that, unlike in the classical preference optimization setting with a human to rank, RLPO preference comes from the TCAV scores (e.g., $TX_{c_{1},t}\succ TX_{c_{2},t}$). We call this notion RLPO-XAI in some ablation studies below. If the generative model at time $t$ is not capable of generating concepts that are in an explainable state, $\max(TS_{c_{1},m,t},TS_{c_{2},m,t})\leq\eta$, we then perform preference update on SD’s weights (more details in Appendix C.5). Following Low-Rank Adaptation (LoRA) (Hu et al., 2021), we only learn auxiliary weights $a$ and $b$ at each time step, and update the weights as $w_{t+1}\leftarrow w_{t}+\lambda ab$.
@@
As the deep RL agent progresses over time, the states become more relevant as it approaches explainable states (Fig. 1), thus the same action yields increasing rewards over time. To accommodate this, with reference to the rewards defined in Section 3.3, we introduce a parameter $\xi$, which starts at 0.1 and incrementally rises up to 1 as the preference threshold, $\eta$, is approached. Different actions may result in different explainable states, reflecting various high-level concepts inherent to $f(\cdot)$. Some actions might take longer to reach an explainable state. Also, it is possible for different actions to lead to the same explainable state. As the goal is to optimize all states to achieve a common target, DQN progressively improves action selection to expedite reaching these states. Thus, deep RL becomes relevant as it optimizes over time to choose the actions that are most likely to reach an explainable state more efficiently.
4 Experiments
To test the effectiveness of our approach, we tested it across multiple models and several classes. We considered three CNN-based classifiers, ResNet-50 (He et al., 2016), GoogleNet (Szegedy et al., 2015), and InceptionV3 (Szegedy et al., 2016), and two transformer-based classifiers, ViT (Dosovitskiy et al., 2020) and Swin (Liu et al., 2021), pre-trained on ImageNet dataset. Unless said otherwise, only GoogleNet results are shown in the main paper. All other model details and results are provided in Appendix C.1 and D.3), respectively. All metrics are defined in Appendix B.1.
4.1 Ablation study: Search Strategies (Why deep RL?)
We chose DQN as our RL algorithm because of its ability to effectiveness traverse through discrete action space Mnih et al. (2015) (20 unique seed prompts). We assess the effectiveness of RL by disabling the preference optimization step. As shown in Table 2, on GoogleNet classifier, compared to $\epsilon$-greedy methods, RL setup exhibits a higher entropy, average normalized count (ANC), and inverse coefficient of variance (ICV) (See Appendix B.1 for definitions), indicating RL’s ability to take diverse actions that results in diverse concepts.
4.2 Ablation study: Scoring Mechanisms (Why TCAV?)
Another important aspect of our setup is the use of TCAV score, an XAI method, to provide preference feedback and calculate rewards. Alternatively, this XAI scoring feedback can also be replaced with human feedback or LLM-based AI feedback. As an additional experiment, to test the effect of human feedback, we conducted human feedback experiment with eight human subjects who provided live human feedback. Further, to evaluate the LLM-based AI feedback, we made use of GPT-4o. More details on the experiment setup and results are provided in Appendix D.2. As shown in Table 4, we concluded that, even though other feedback techniques can be used, XAI-based feedback is best for generating concepts that are important to model with high speed and low computation cost. Though human and AI (GPT4o) are good at correlating semantics, by only looking at test images and concepts instead of model activations, they are not able to provide model specific explanations. Also, human experiments tend to be expensive.
Table 1: Search strategy ablation. We see that RL, compared to $\epsilon$-greedy search, is the best strategy to efficiently explore the search space with high entropy, high average normalized count (ANC) per action, and high inverse coefficient of variance (ICV).
Method Entropy ($\uparrow$) ANC ($\uparrow$) ICV ($\uparrow$)
RL (Ours) 2.80 0.43 2.17
0.25 Greedy 2.40 0.21 1.04
0.5 Greedy 1.95 0.15 0.59
0.75 Greedy 1.85 0.15 0.56
Table 2: Exploration Gap (EG) and Odds calculated based on the responses for ours and retrieval based method, respectively, from the human survey (Appendix D.6).
Laymen Expert
(n=260) (n=240)
EG (Retrieval) 6.54% 10.45%
EG (Ours) 91.54% 65.45%
Odds (Retrieval) 14.29 8.57
Odds (Ours) 0.09 0.53
Table 3: Scoring mechanisms ablation. We see that RLPO with Explainable AI feedback (RLPO-XAIF), in this case TCAV, is a better choice than RLPO with human feedback (RLPO-HF) and AI feedback (RLPO-AIF).
Method Class-based Model-specific Feedback Execution
Explanations Explanations Cost* Time (↓)
RLPO-HF NIL 180 $\pm$ 30s
RLPO-AIF >10 GB 72 $\pm$ 1.2s
RLPO-XAIF <1 GB 56 $\pm$ 0.7s
* Feedback cost refers to the storage or memory requirements for feedback processing.
Table 4: Novel concepts: $TS_{c,m}$ (TCAV score), CS (Cosine similarity), ED (Euclidean distance), RCS, and RED (CS and ED with ResNet50 embedding)
Model Concepts $TS_{c,m}(\uparrow)$ CS ($\downarrow$) ED ($\uparrow$) RCS ($\downarrow$) RED ($\uparrow$)
EAC (Sun et al., 2024) C $1.0$ $0.76\pm 0.03$ $7.21\pm 0.63$ $0.67\pm 0.14$ $6.34\pm 2.16$
Lens (Fel et al., 2024) C1 $1.0$ $0.77\pm 0.02$ $7.17\pm 0.34$ $0.50\pm 0.18$ $9.70\pm 3.20$
C2 $1.0$ $0.72\pm 0.04$ $8.02\pm 0.87$ $0.42\pm 0.10$ $10.90\pm 2.80$
C3 $1.0$ $0.69\pm 0.05$ $8.45\pm 0.96$ $0.45\pm 0.05$ $11.03\pm 2.17$
CRAFT (Fel et al., 2023) C1 $1.0$ $0.76\pm 0.04$ $7.37\pm 0.62$ $0.57\pm 0.16$ $8.80\pm 3.20$
C2 $1.0$ $0.72\pm 0.02$ $8.25\pm 0.39$ $0.50\pm 1.90$ $9.90\pm 3.40$
C3 $1.0$ $0.73\pm 0.04$ $7.98\pm 0.79$ $0.44\pm 0.07$ $10.80\pm 1.90$
RLPO (Ours) C1 $1.0$ $0.52\pm 0.04$ $10.48\pm 0.50$ $0.04\pm 0.01$ $16.80\pm 1.40$
C2 $1.0$ $0.49\pm 0.02$ $10.65\pm 0.20$ $0.02\pm 0.02$ $17.20\pm 0.80$
C3 $1.0$ $0.49\pm 0.02$ $10.74\pm 0.30$ $0.03\pm 0.01$ $17.60\pm 4.40$
4.3 What kind of concepts can RLPO generate?
Novel concepts. As illustrated in Fig. 4, we observed that the RLPO can generate concepts that a human would not typically think of but leads activations of the DNN to trigger. To validate this hypothesis, we conducted a survey to see if humans can think of these generated concepts as important for the neural network to understand a certain class (Table 2, The Exploration Gap (EG) quantifies the proportion of missed optimal actions, defined as $1-\text{Accuracy}$, highlighting how humans frequently miss the most optimal actions when presented with generated concepts.). As detailed in Appendix D.6, we presented a random class image followed by two concepts, one generated by our method and another from a previous retrieval based method (Fel et al., 2024; 2023). We presented choices from the concepts with high score from both methods and we discovered that while most participants could recognize retrieval-based concepts, only those with domain-specific knowledge could identify generated concepts. This indicates that most people can only identify concepts from a small subset of what $f(.)$ learns during training. Intuitively, when we retrieve concepts from the test class, they tend to be similar to the test images.
Figure 4: Samples of concepts (with different TCAV scores, not shown here) generated by different methods. Observe that RLPO generates diverse images, not just patches from the test images.
We also compare the generated concepts qualitatively and quantitatively. Fig. 4 shows the diversity of concepts generated by our method and other retrieval based methods. Additionally, we also verify this by computing the vector similarity of the CLIP and ResNET50 embeddings between $X_{c}$ and $X_{m}$ for multiple retrieval based methods. As highlighted in Table 4, we observe that retrieval based methods tend to have high cosine similarity between extracted concepts and test images, making them less useful as abstract concepts (e.g., to explain the zebra class, a patch of zebra as a concept is less useful compared to stripes concept).
Abstract concepts. In Fig. 5, we observe the progression of output concepts generated by the SD when RLPO is applied for the seed prompt “zoo,” of tiger class. These abstractions hint us about what the model prefers when it is looking for tiger, starting from a four-legged orange furred animal, to black and white stripes with orange furred animal, to black and white stripes with orange furred and whiskers. We obtain concepts with various abstractions by changing $\eta$ (Currently, it is not possible for our method to decide $\eta$ to get a particular level of abstraction).
Figure 5: Different levels of abstractions for the “Tiger” class on GoogleNet classifier. The generated image begins as a random zoo image and gradually transition to images with tiger features. We also see that the model’s prediction changes over time, starting from a random classification of oxcart to classifying the generated concept as tiger itself.
Multiple concepts. Because RLPO algorithm explores various explainable states, we can obtain multiples concepts with varying level of importance. Fig. 6 shows the top three class-level concepts identified by our method for the “zebra” class for the GoogleNet classifier. We see that, each concept set has a different TCAV score associated with them indicating their importance.
4.4 Are the generated concepts reliable?
Figure 7: Comparison of concepts identified by different methods. RLPO can show the correspondences between test image and different concepts.
After generating the concepts, next step is to identify what those concepts signify. To locate where in the class images generated concepts correspond, we made use of CLIPSeg (Lüddecke & Ecker, 2022), a transformer-based segmentation model which takes in concept images as prompts, $X_{c}$, and highlights in a test image, $x\in X_{m}$, which part resembles the input prompt as a heat map. More details on this is available in Appendix D.3.3. As shown in Fig. 6, class image on left highlights the top 3 identified concepts by RLPO. We also compare the output generated by other popular XAI techniques such as LIME and GradCam with ones generated by RLPO. As shown in Fig. 7, we can see that other methods just explains where the model is looking at whereas our approach also explains what type of features is the model focuses on.
After finding the relationship between generated concepts and input images, we need to validate the importance of the identified concepts. To that end, we applied c-deletion, a commonly used validation method in XAI, to the class images for each identified concept. We gradually deleted concept segments based on the segmentation heat map obtained from ClipSeg. The results for the c-deletion are shown in the Fig. 9. We see the area under curve is the highest for the most important concept “stripes” and the lowest the least important concept “mud,” indicating the order of importance of each concept. More examples on the c-deletion are in Appendix D.3.4.
Figure 8: C-deletion. Removing concepts over time to measure the reliability. The colored numbers indicate the area under the curve (the lower the better as it indicates important concepts are removed sooner.)
Figure 9: How some eight concepts are shifted from pre (blue) to post (orange) fine-tuning of ResNet50 model on Blue Jay class.
Figure 10: We use sentiment analysis in NLP to show the generalizability of RLPO. Here, concepts tend to be synonyms and the numbers indicate the TCAV scores. Generated concepts explain why a given text is classified as a positive sentiment.
4.5 Applications and generalizability
We show how RLPO can be used as a diagnostic tool for the engineers. As a specific application, we see what concepts are removed and added, as well as how the concept importance changes when we fine-tune ResNet50 model on ImageNet to improve accuracy (Fig. 9). More details about the experiment is present in Appendix D.5.
To demonstrate the generalizability of the proposed algorithm, we extended RLPO to generate words in sentiment analysis in NLP. We made use of Mistral-7B Instruct model to generate synonyms of seed prompts and optimized the language model based on preferences from TextCNN model pre-trained on IMDB sentiment dataset. Fig. 10 highlights relevant words with their importance score in the input. More details about the experiment is present in Appendix D.4.
5 Limitations and Conclusions
The process of navigating an infinitely large concept space and generating explainable concepts from textual inputs presents several challenges, particularly when dealing with complex, high-level concepts. We showed how deep RL can guide SD efficiently to navigate this space. However, RLPO also suffers from several limitations. First, this analysis cannot be performed real-time since generating images from SD, learning the DQN, and fine-tuning the SD with preferences takes some time. Also, the concepts that our algorithm generates can be diverse as it tries to reveal the concepts inherent to the NUT, making it less domain-specific (e.g., for a medical application, there is a chance it might generate non-medical images if the NUT activations get excited for non-medical data). As a future extension, we hope to input preferences from both TCAV and application experts while optimizing, making generated explanations even more aligned to specific applications. Despite the challenges, our results show how to leverage the strengths of visual representations and adaptive learning to provide intuitive and effective solutions for understanding complex, high-level concepts in black-box neural networks.
Appendix
Appendix A Related Work
Use of VLM in explaination: Recent advancements in Vision-Language models (VLMs) have open the doors for the use of VLMs in multiple domains, mainly because of their ability to generalize over large amount of data, they can be leveraged to obtain useful information. Work by Sun et al. (2024) present a novel method combining the Segment Anything Model (SAM) with concept-based explanations, called Explain Any Concept (EAC). This method uses SAM for precise instance segmentation to automatically extract concept sets from images, then it employs a lightweight surrogate model to efficiently explain decision made by any neural network based on extracted concepts. Another work by Yan et al. Yan et al. (2023) introduced Learning Concise and Descriptive Attributes (LCDA), which leverages Large Language Models (LLMs) to query a set of attributes describing each class and then use that information with vision-language models to classify images. They highlight in their paper that with a concise set of attributes, they can improve the classifier’s performance and also increase interpretability and interactivity for end user.
Appendix B Definition, Theorems, and Proves
Comparative Overlap of Human-Interpretable and Generative Model Concepts in Neural Understanding Tasks (NUT). We formalize this with reference to Fig. 2.
@@ Appendix B Definition, Theorems, and Proves
Proof of Theorem 1.
Definition 3.
Let $\mathcal{C}_{H}$, $\mathcal{C}_{R}$, and $\mathcal{C}_{G}$ denote the sets representing human-interpretable concepts, retrieved concepts, and concepts generated by a generative model, respectively. We define the relationships between these sets as follows:
$\mathcal{C}_{H}\subseteq\mathcal{C}_{G}\quad\text{and}\quad\mathcal{C}_{R}\subseteq\mathcal{C}_{G}.$
Property 1.
For any set $\mathcal{C}_{i}$, where $i\in\{H,R,G\}$, it holds that:
$\emptyset\subseteq(\mathcal{C}_{i}\cap\mathcal{C}_{N})\subseteq(\mathcal{C}_{i}\cup\mathcal{C}_{N}),$
where $\mathcal{C}_{N}$ represents the set of concepts learned by the NUT.
For any two sets $\mathcal{A}$ and $\mathcal{B}$, the size of their intersection $|\mathcal{A}\cap\mathcal{B}|$ is non-negative since it represents the number of elements common to both sets. Thus, we have:
$|\mathcal{C}_{H}\cap\mathcal{C}_{N}|\geq 0\quad\text{and}\quad|\mathcal{C}_{R}\cap\mathcal{C}_{N}|\geq 0.$ (4)
Given Definition 3 and Property 1, we assume the following subset relationships between the sets:
$|\mathcal{C}_{H}\subseteq\mathcal{C}_{G}|\quad\text{and}\quad|\mathcal{C}_{R}\subseteq\mathcal{C}_{G}|.$
Case 1. Since $\mathcal{C}_{H}\subseteq\mathcal{C}_{G}$, any element $x\in\mathcal{C}_{H}$ is also in $\mathcal{C}_{G}$. Therefore, any element $x\in\mathcal{C}_{H}\cap\mathcal{C}_{N}$ is also in $\mathcal{C}_{G}\cap\mathcal{C}_{N}$. Hence,
$\mathcal{C}_{H}\cap\mathcal{C}_{N}\subseteq\mathcal{C}_{G}\cap\mathcal{C}_{N}.$
Case 2. Similarly, since $\mathcal{C}_{R}\subseteq\mathcal{C}_{G}$, any element $x\in\mathcal{C}_{R}$ is also in $\mathcal{C}_{G}$. Therefore, any element $x\in\mathcal{C}_{R}\cap\mathcal{C}_{N}$ is also in $\mathcal{C}_{G}\cap\mathcal{C}_{N}$. Hence,
$\mathcal{C}_{R}\cap\mathcal{C}_{N}\subseteq\mathcal{C}_{G}\cap\mathcal{C}_{N}.$
From Case 1 and Case 2, since $\mathcal{C}_{H}\cap\mathcal{C}_{N}$ and $\mathcal{C}_{R}\cap\mathcal{C}_{N}$ are subsets of $\mathcal{C}_{G}\cap\mathcal{C}_{N}$, it follows that:
$|\mathcal{C}_{H}\cap\mathcal{C}_{N}|\leq|\mathcal{C}_{G}\cap\mathcal{C}_{N}|$ (5)
and
$|\mathcal{C}_{R}\cap\mathcal{C}_{N}|\leq|\mathcal{C}_{G}\cap\mathcal{C}_{N}|$ (6)
Combining Eqs. 5 and 6 with the non-negativity established in Eq 4, we have:
$|\mathcal{C}_{G}\cap\mathcal{C}_{N}|\geq|\mathcal{C}_{H}\cap\mathcal{C}_{N}|\geq 0$ (7)
and
$|\mathcal{C}_{G}\cap\mathcal{C}_{N}|\geq|\mathcal{C}_{R}\cap\mathcal{C}_{N}|\geq 0.$ (8)
The Effects of DQN-DPO-based Concept Space Traversal. We now formalize what concepts the DQN has learned, with reference to Fig. 1.
@@
Property 2.
As $\xi$ increases, the reward function proportionally amplifies, particularly enhancing the significance of outcomes near $t_{\eta}$.
Proof of Theorem 2.
WLOG, let the TCAV score, $S_{c,m,t}$, for concept $c$ and class $m$ at time $t$ be $S_{t}$. The reward function is defined as (Section. 3.3),
$R(t,a)=K\cdot S_{t}\cdot f(t),$ (9)
for a constant $K$ and a factor,
$f(t)=\begin{cases}\xi\cdot t&\text{if }t\leq t_{\eta},\\ \xi_{0}&\text{otherwise},\end{cases}$ (10)
for positive parameters $\xi$ and $\xi_{0}$.
Case 1: Considering the difference in reward function at time $t$ when $t\leq t_{\eta}$,
$R(t+1,a)-R(t,a)$ $=K\cdot S_{t+1}\cdot f(t+1)-K\cdot S_{t}\cdot f(t)$ (11)
$=K\cdot S_{t+1}\cdot\xi\cdot(t+1)-K\cdot S_{t}\cdot\xi\cdot t$
$=K\cdot\xi\cdot(S_{t+1}\cdot(t+1)-S_{t}\cdot t)$
$=K\cdot\xi\cdot(t(S_{t+1}-S_{t})+S_{t+1})$
$\text{From }(S_{t+1}-S_{t})=\frac{h(t+1)-h(t)}{(t+1)-t})=h^{\prime}(t),$ (12)
$R(t+1,a)-R(t,a)$ $=K\cdot\xi\cdot(t\cdot h^{\prime}(t)+S_{t+1}).$ (13)
Since,
  1. 1.
    $S_{t}$ is monotonically increasing for $t\leq t_{\eta}\implies h^{\prime}(t)>0$ and
  2. 2.
    $S_{t}\in[0,1]$,
$R(t+1,a)-R(t,a)\geq 0.$ (14)
Case 2: Considering the same difference in rewards for $t\geq t_{\eta}$.
$R(t+1,a)-R(t,a)$ $=K\cdot S_{t+1}\cdot f(t+1)-K\cdot S_{t}\cdot f(t)$ (15)
$=K\cdot S_{t+1}\cdot\xi-K\cdot S_{t}\cdot\xi_{0}$
$=K\cdot\xi_{0}\cdot(S_{t+1}-S_{t})$
Given that $S_{t}$ and $S_{t_{+}1}$ are both outcomes generated from a generative model fine-tuned for a particular concept, $S_{t+1}-S_{t}\approx 0$ in response to the same action. Hence,
$R(t+1,a)-R(t,a)\approx 0.$ (16)